Circular No. 6068

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

POSSIBLE GRAVITATIONAL MICROLENSING EVENT The MACHO collaboration (A. Alcock, T. Axelrod, R. Allsman, D.

Bennett, K. Cook, K. Freeman, K. Griest, S. Marshall, B. Peterson, M. Pratt, P. Quinn, D. Reynolds, A. Rodgers, C. Stubbs, and W. Sutherland; Lawrence Livermore National Laboratory, Center for Particle Astrophysics, and Mount Stromlo Observatory) report: “We have discovered a possible gravitational microlensing event in progress toward the galactic bulge. The star in question is located at α = 17^{h}59^{m}49^{s}.6, δ = −28^{o}10′56′′ (equinox 2000.0); it was constant at R = 17.8, V = 18.8 during 1993 and mid-1994, and has brightened by 0.85 mag during the week ending Aug. 31; the color appears unchanged. Magnitude estimates from the 1.27-m Mt. Stromlo telescope (+ CCD): Aug. 29.39 UT. R = 17.24; 30.43, 17.08; 31.43, 16.97 (errors typically 0.05 mag). If this event is due to gravitational microlensing, a preliminary fit predicts that the star should reach a peak of R < 16.6 between Sept. 3 and 9 and then decline symmetrically. As the light curve is currently incomplete, it is not possible to confirm that this is definitely a microlensing event, so it should be emphasized that the microlensing interpretation is still tentative. Accurate multicolor photometry and especially spectroscopy of this star at regular intervals during the next 14 days will be very valuable to test the microlensing interpretation. Light curves and a finding chart are available by anonymous FTP to lensing.physics.ucsb.edu in the directory macho/alert/event2. This directory will be periodically updated with new light curves and microlensing fits as the data come in. Potential observers are requested to contact D. Bennett (telephone USA 510-423-0656; e-mail bennett@sunlight.llnl.gov) to coordinate observations.”

SUPERNOVA 1994X IN ANONYMOUS GALAXY S. Benetti, European Southern Observatory (ESO), writes: “I obtained

a spectrum (450–710 nm, resolution 1.1 nm) of SN 1994X (cf. IAUC 6056) on Aug. 29.35 UT with the ESO 2.2-m telescope (+ EFOSC2) at La Silla. The full reduction shows the broad features of a typical type-Ia supernova spectrum ∼ 3 weeks after maximum. By comparing the wavelength of the sharper emissions with those of other well-studied type-Ia supernovae at the same phase, a redshift of z ∼ 0.056 is deduced for the parent galaxy. Assuming this redshift, an expansion velocity of ∼ 8900 km/s is derived from the minimum of the Si II 635.5-nm multiplet; this is a typical expansion velocity for a type-Ia supernova three weeks after maximum.”

1994 September 1 Daniel W. E. Green